# Statistical Advantage of Softmax Attention: Insights from Single-Location Regression

### v1

@@ 1 Introduction
For these reasons, numerous alternatives have been proposed. Notably, kernelized attention approximates the softmax function with kernel feature maps (Wang et al., 2020; Luo et al., 2021; Choromanski et al., 2021; Qin et al., 2022), achieving linear complexity in sequence length. In parallel, state-space models (Peng et al., 2023; Poli et al., 2023; Gu and Dao, 2024) introduce linear recurrent dynamics with gating to tackle long-context information retrieval. On the other hand, while less studied empirically, linear attention, which is the linearization of softmax around the origin, has been extensively studied in theoretical works (e.g. Ahn et al., 2023; von Oswald et al., 2023; Bai et al., 2023; Mahankali et al., 2023; Zhang et al., 2024; Lu et al., 2025; Zhang et al., 2025).
The comparison of softmax with these alternatives has been investigated in particular in the framework of synthetic in-context retrieval tasks, such as Needle-in-a-Haystack (NIAH), where the model must retrieve a single fact hidden in a long block of irrelevant text (Kamradt, 2023; Machlab and Battle, 2024). Overall, the comparison is nuanced: while some alternatives, in particular SSMs, match or even outperform softmax in some of these synthetic tasks, they fail to do so in slight variations thereof or in actual language settings (Arora et al., 2024; Hsieh et al., 2024; Shen et al., 2024). All in all, the practical dominance of softmax in attention is still not fully understood.
@@
In this paper, we take a more principled approach to the question, with the following contributions. Code to reproduce our simulations is provided.
  • We propose a mathematical formalization of tasks where the output depends on a single input token as the single-location regression model (Section 3). This model subsumes previous theoretical studies (Marion et al., 2025), notably by introducing random sequence lengths, and gives a formalization of information retrieval tasks such as NIAH.
@@
  • We prove a gap in approximation performance between linear and softmax attention (Section 4), with the latter reaching Bayes performance. We argue that this analysis captures the performance efficiently attainable with one-pass stochastic gradient descent. This gap arises from both the exponential nonlinearity and the normalization in softmax, as illustrated by our comparison with kernelized and element-wise attention (Fig. 1).
@@ 2 Further Related work
While SSMs outperform attention-based models in AR thanks to their capacity to handle long contexts, this success is brittle since they lag behind Transformers on the slightly more involved MQAR task (Arora et al., 2024; Aksenov et al., 2024; Chou et al., 2024; Wang et al., 2025). Arora et al. (2024) propose a theoretical explanation based on the lack of expressivity of convolution-gating models. Transformers also outperform SSMs and kernelized attention in (a generalization of) NIAH (Hsieh et al., 2024; Shen et al., 2024). In this work, we take a principled approach to understanding the benefits of softmax, by identifying an information retrieval framework amenable to theoretical study, and going beyond expressivity results to statistical and computational advantages.
@@
While our data model is inspired by Marion et al. (2025), it departs significantly by generalizing to variable sequence length, introducing a generic weighting mechanism to encode single location, and renormalizing to obtain proper high-dimensional limits, thereby allowing theoretical analysis of the softmax nonlinearity. The data model belongs among the sequence multi-index models as introduced in Cui (2025). Special cases of single-index models we studied in Cui et al. (2024); Arnaboldi et al. (2025). Our model corresponds to a sequence two-index case, for which the Bayes-optimal estimator was studied in Troiani et al. (2025). A related classification problem where outputs only depend on a few tokens was concurrently studied in Barnfield et al. (2025) with analysis of steps-wise gradient descent training.
@@ 3.2 Probabilistic data model for SLR
$k^{*}\sim\mathcal{N}(0,I_{D})\ ,\quad v^{*}\sim\mathcal{N}(0,I_{D})$ (1)
@@ 4 Comparison of softmax with alternatives in population risk
In this section, we compare the expressivity of the softmax attention with alternatives on our task, by theoretically assessing the Bayes risk and the minimal population risk for various activation functions. We show that the softmax reaches the Bayes performance. We combine analytical arguments with numerical evidence to support the claim that this approach also captures the performance that is efficiently achievable by running one-pass SGD, which directly minimizes the population risk from random initialization. All proofs of the results in this section are deferred to Appendix A.
@@ 4.1 Bayes risk and optimality of softmax
We turn our attention to the best reachable theoretical risk over the class of estimators $\mathcal{F}_{\sigma}$ when $\sigma$ is the softmax function. Surprisingly, this estimator reaches the Bayes performance as shown next.
@@
Note that the requirement on $g_{\nu}$ defining the distribution of $X$ holds in particular for the spiked-SLR and the max-SLR. In consequence, the softmax architectures $\mathcal{F}_{\rm softmax}$ reach the Bayes performance in both settings, and furthermore the proof reveals that the minimum is reached for $k=c_{\nu}k^{*}$ and $v=v^{*}$, corresponding to recovery of the hidden directions. In statistical physics terminology, the softmax attention is said to satisfy the Nishimori condition, as detailed in the proof. This is to be contrasted with the performance of other activation functions, which is characterized next.
@@ 4.2 Expression of the population risk for arbitrary activation functions
The population risk of the attention $(k,v)\mapsto\mathcal{E}_{\sigma}(k,v)$ can be reparameterized as a function of the following 7 variables, referred to as order parameters:
@@
The first two order parameters, $m_{kk^{*}}$ and $m_{vv^{*}}$, quantify the recovery of the hidden directions, while $q_{kk}$ and $q_{vv}$ are the squared norm of the parameters $k$ and $v$. Finally, the cross-correlation terms $m_{kv^{*}}$, $m_{vk^{*}}$, and $q_{vk}$ are nuisance parameters. We let $\tilde{\mathcal{E}}_{\sigma}:\mathbb{R}^{7}\to\mathbb{R}$ be the reparameterized risk depending on the order parameters.
@@
This simplification is supported by the following observations pertaining to the landscape of the population risk $\mathcal{E}_{\sigma}$. In practice, we do not have access to minimizers of the risk, but can instead optimize parameters $k$ and $v$ by one-pass SGD on $\mathcal{E}_{\sigma}$, which corresponds on average to running gradient descent (GD) on the reparameterized risk $\tilde{\mathcal{E}}_{\sigma}$. In the high-dimensional limit $D\to\infty$, random Gaussian initialization of the parameters lands on $\mathcal{M}$ because the cross-correlations vanish. Then, the manifold is invariant by GD, in the sense that iterates initialized on the manifold remain on it (see Appendix A.2 for details on this statement and the following). At finite $D$, random initialization lands in a neighborhood of $\mathcal{M}$. In this case, we numerically check that $\mathcal{M}$ is stable in the sense that parameters stay close throughout GD. For linear attention, we also show that the risk has only one (local) minimizer on $\mathcal{M}$, which is also a local minimizer of the risk over the whole parameter space, while for softmax attention we know that a global minimizer is on $\mathcal{M}$. Taken together, these facts support that GD on $\tilde{\mathcal{E}}_{\sigma}$ and thus one-pass SGD on $\mathcal{E}_{\sigma}$ converges to a minimizer of the risk on $\mathcal{M}$. A formal proof is delicate, as the landscape features other local minima outside of the manifold.
With a small abuse of notation, we let $\tilde{\mathcal{E}}_{\sigma}(m_{kk^{*}},m_{vv^{*}},R_{kk},R_{vv})$ be the reparameterized risk restricted to $\mathcal{M}$, with $R_{kk}^{2}=q_{kk}-m_{kk^{*}}^{2}$ and $R_{vv}^{2}=q_{vv}-m_{vv^{*}}^{2}$ the orthogonal components. It turns out that this object is amenable to insightful theoretical analysis. We first write it explicitly.
@@ 4.3 Comparison between linear and softmax attentions
Letting $\mathsf{E}_{\sigma}=\min_{\mathcal{M}}\mathcal{E}_{\sigma}=\min_{\mathbb{R}^{4}}\tilde{\mathcal{E}}_{\sigma}$ be the minimum of the population risk, we now compare the optimal population risk $\mathsf{E}_{\mathrm{lin}}$ for linear attention $\sigma(\chi)_{\ell}=1+\chi_{\ell}$ to the one for softmax $\mathsf{E}_{\mathrm{softmax}}=\mathcal{E}_{\mathrm{Bayes}}$, encompassing asymptotic regimes of strong signals $\nu\to\infty$ or long sequences $L\to\infty$.
@@
$\mathsf{E}_{\mathrm{lin}}=1-\frac{L+\nu(L-1)}{L^{2}+\nu(L-1)}{=}\frac{L}{L-1}\frac{1}{\nu}+o_{\nu\to\infty}(1)\quad\text{while }\quad\mathsf{E}_{\mathrm{softmax}}=e^{-c_{L}\nu+o_{\nu\to\infty}(\nu)}$ (12)
@@
Figure 1: Minimal population risk $\mathsf{E}_{\sigma}$ over $\mathcal{F}_{\sigma}$ for different attention activations $\sigma$ (colors), compared to the Bayes risk $\mathcal{E}_{\mathrm{Bayes}}$ (5) (black), for the two tasks spiked-SLR (top) and max-SLR (bottom). Softmax is the only one achieving the Bayes risk. The markers on the lines are for readability only. Population risks are computed via numerical optimization of (9), starting from a random initialization. In all cases, we found that $R_{kk}=R_{vv}=0$ was optimal, i.e. $k$ exactly aligns with $k^{*}$ and $v$ with $v^{*}$.
@@ 5 Performance of the attention at finite sample complexity
So far, our analysis has focused on the properties of the minimizer of the population risk. In practice, however, we only have access to a finite dataset $\mathcal{D}=\left\{(X_{\mu},y_{\mu})\in\mathbb{R}^{L(\mu)\times(D+1)}:\mu\in[N]\right\}$, and must therefore rely on an empirical estimator. In this section, we consider the regularized empirical risk defined for some regularization strengths $r_{k}\geq 0$ and $r_{v}\geq 0$ by
@@
We investigate the high-dimensional proportional limit where $N,D\to\infty$ at finite $\nicefrac{{N}}{{D}}\to\alpha=\Theta(1)$. Observe that letting $\alpha\to\infty$ recovers the population risk discussed in Section 4. For simplicity, we consider the noiseless task $\Delta=0$ in the following.
While in Section 4 the benchmark was the Bayes risk, at finite sample complexity it is natural to consider instead the best achievable test risk conditionally on a finite batch of data $\mathcal{D}$, also known as the Bayes-optimal risk. In the proportional high-dimensional limit, the Bayes-optimal performance for the SLR task can be computed by extending the results of Troiani et al. (2025) to random sequence lengths $L\sim P_{L}$ and arbitrary $g_{\nu}$. Although this result is of independent interest, for conciseness we refer the reader to Appendix C for the details. The outcome is that the Bayes-optimal performance presents a rich phenomenology, with a hard phase where the best-known first-order method fails to achieve the information-theoretical performance, see Appendix B for a full discussion.
@@
As in the population result of Proposition 4.3, the high-dimensional analysis shows that the risk depends only on a few order parameters. Although the resulting expressions are cumbersome, it is important to emphasize that these quantities are deterministic and independent of the diverging dimensions $N$ and $D$, in contrast to the original risk (15). Thus the result provides an implementable formula for the performance of ERMs. In practice, the minimum over the set $\mathcal{S}$ is computed by running the fixed point method from several initializations, typically random (so-called uninformed) and informed ones corresponding to (partial) alignment of $k$ and $v$ with $k^{*}$ and $v^{*}$.
@@
Figure 2: Test error of the trained attention (linear vs. softmax) across different tasks and signal strengths $\nu$, for $L=3$. Linear attention is shown in red and softmax in blue. Solid lines indicate $\mathsf{E}_{\sigma}(\alpha)$ at finite $\alpha$ (Result 5.1), while markers represent the test error of an ERM obtained via a local optimization method with $\sqrt{ND}=10^{4}$. The regularizations $r_{k}$ and $r_{v}$ are tuned by grid search to minimize the test error. Dotted and dashed lines correspond to the value of $\mathsf{E}_{\sigma}$ in the infinite-$\alpha$ limit (see closed-formed formulas in Proposition 4.1 for softmax and Appendix A.2.7 for linear). The Bayes-optimal risk is shown in black (see Appendix B for a discussion on its discontinuity). Appendices D-E include more experimental details and results.
@@
We conclude by a few computational remarks. Because the empirical risk (15) is non-convex, it is not guaranteed a priori that optimization algorithms can find global minima. To assess the effect of non-convexity, we rely on numerical simulations shown in Fig. 2. More precisely, we compare the prediction of Result 5.1 (here, uninformed and informed initializations give the same result) to the outcome of running a local optimization algorithm (specifically, a quasi-Newton method) on the risk (15), starting from a random initialization, for finite but large $N$ and $D$. We first note that the agreement between both is excellent. This suggests that potential bad local minima in the optimization landscape are avoided, at least for the appropriate regularization strength. Analyzing this landscape is an interesting open question. Furthermore, the softmax attention has lower error than the linear attention across all the tested hyperparameters, which shows that the benefits of softmax extend beyond the population risk analysis of Section 4 to statistical and computational advantages on the empirical risk. We finally note that in this case the softmax is no longer Bayes-optimal but the gap to the Bayes risk closes as $\alpha$ grows, as expected from our analysis in Section 4.

### v2

@@ 1 Introduction
For these reasons, numerous alternatives have been proposed. Notably, kernelized attention approximates the softmax function with kernel feature maps (Wang et al., 2020; Luo et al., 2021; Choromanski et al., 2021; Qin et al., 2022), achieving linear complexity in sequence length. In parallel, state-space models (SSMs, Peng et al., 2023; Poli et al., 2023; Gu and Dao, 2024) introduce linear recurrent dynamics with gating to tackle long-context information retrieval. On the other hand, while less studied empirically, linear attention, which is the element-wise linearization of softmax around the origin, has been extensively studied in theoretical works (e.g. Ahn et al., 2023; von Oswald et al., 2023; Bai et al., 2023; Mahankali et al., ; Zhang et al., 2024; Lu et al., 2025; Zhang et al., 2025).
Despite substantial research on alternative attention mechanisms, softmax attention remains the dominant architecture in large-scale language models. A leading hypothesis for this empirical success is its superior performance on retrieval tasks, a view supported by a growing body of work (Arora et al., 2024; Aksenov et al., 2024; Chou et al., 2024; Hsieh et al., 2024; Shen et al., 2024; Wang et al., 2025). While we defer a detailed discussion to Section 2, we briefly summarize the representative study of Shen et al. (2024). These authors characterize optimal scaling laws for language models equipped with kernelized attention and state-space models (SSMs), and then train softmax-attention, kernelized-attention, and SSM-based models with matched parameter budgets ranging from 7070M to 77B. Their downstream evaluation, reproduced for convenience in Figure 1, reveals a striking pattern: although optimally-scaled SSMs and kernelized-attention models achieve competitive performance on linguistic proficiency benchmarks, they consistently underperform softmax attention on retrieval tasks.
Yet the underlying reasons for the advantage of softmax in retrieval tasks remain poorly understood. This gap limits our ability to reason about the tradeoffs and fundamental constraints of different model architectures.
(a) Retrieval tasks (Needle-in-a-Haystack, Kamradt, 2023; Machlab and Battle, 2024). Colors correspond to various context lengths (longer context makes the task harder). The “average” curve is the average over the three tested context lengths.
(b) Linguistic proficiency tasks (accuracy on Common Sense Reasoning and SCROLLS, perplexity on WIKITEXT-2 and LAMBADA datasets). All metrics are normalized so that higher is better, and 11 (resp. 0) corresponds to the best (resp. worse) performance over all tested configurations.
Figure 1: (Shen et al., 2024, Tables 4 and 11): Comparison of softmax Transformer (LLaMA, bolded line) with kernelized attention (TNL, cos2) and state-space models (HGRN2), as a function of the model size, and for various tasks (retrieval tasks on the left, linguistic proficiency on the right). All architectures have similar performance for the linguistic proficiency tasks, whereas in retrieval tasks the softmax attention systematically outperforms alternatives.
@@
In this paper, we take a principled approach to the question, with the following contributions. Code to reproduce our simulations is provided.
  • We propose a mathematical formalization of tasks where the output depends on a single input token as the single-location regression model (Section 3). This model subsumes previous theoretical studies (Marion et al., 2025), notably by introducing random sequence lengths, and gives a formalization of information retrieval tasks.
@@
  • We prove a gap in approximation performance between linear and softmax attention (Section 4), with the latter reaching Bayes error. We argue that this analysis captures the performance efficiently attainable with one-pass stochastic gradient descent. This gap arises from both the exponential nonlinearity and the normalization in softmax, as illustrated by our comparison with kernelized and element-wise attention (Fig. 2).
@@ 2 Further Related work
While SSMs outperform attention-based models in the specific synthetic task of AR thanks to their capacity to handle long contexts, this success is brittle since they lag behind Transformers on the slightly more involved MQAR task (Arora et al., 2024; Aksenov et al., 2024; Chou et al., 2024; Wang et al., 2025). Arora et al. (2024) propose a theoretical explanation based on the lack of expressivity of convolution-gating models. Transformers also outperform SSMs and kernelized attention in retrieval tasks based on linguistic data (Hsieh et al., 2024; Shen et al., 2024). In this work, we take a principled approach to understanding the benefits of softmax, by identifying an information retrieval framework amenable to theoretical study, and going beyond expressivity results to statistical and computational advantages.
@@
While our data model is inspired by Marion et al. (2025), it departs significantly by generalizing to variable sequence length, introducing a generic weighting mechanism to encode single location, and renormalizing to obtain proper high-dimensional limits, thereby allowing theoretical analysis of the softmax nonlinearity. The case of diverging signal strength was concurrently studied in Dohmatob (2025). The data model belongs among the sequence multi-index models as introduced in Cui (2025). Special cases of single-index models we studied in Cui et al. (2024); Arnaboldi et al. (2025). Our model corresponds to a sequence two-index case, for which the Bayes-optimal estimator was studied in Troiani et al. (2025). A related classification problem where outputs only depend on a few tokens was concurrently studied in Barnfield et al. (2026) with analysis of steps-wise gradient descent training.
Last, the concurent work Dragutinović et al. (2025) also discusses advantages of softmax over linear attention, but on a different model (in context classification) and proof techniques.
@@ 3.2 Probabilistic data model for SLR
$k^{*}\sim\mathcal{N}(0,I_{D})\ ,\quad v^{*}\sim\mathcal{N}(0,I_{D})\ .$ (1)
@@ 3.3 Learning with attention
Table 1: Terminology and notations for the risks. In the text, we use the terms “risk” and “error” in an interchangeable manner.
$\mathcal{E}_{\mathrm{Bayes}}$ Bayes risk (in population)
$\mathcal{E}_{\sigma}(k,v)$ population risk of the attention over $(\mathbb{R}^{D})^{2}$
$\tilde{\mathcal{E}}_{\sigma}$ population risk parameterized over $\mathbb{R}^{7}$ or $\mathbb{R}^{4}$
$\mathsf{E}_{\sigma}$ minimal population risk
$\mathcal{E}_{\mathrm{BO}}(\alpha)$ Bayes-optimal risk (empirical)
$\mathcal{L}(k,v)$ regularized empirical risk (loss)
[…]
$\mathsf{E}_{\sigma}(\alpha)$ minimal test risk
@@ 4 Comparison of softmax with alternatives in population risk
In this section, we compare the expressivity of the softmax attention with alternatives on our task, by theoretically assessing the Bayes risk and the minimal population risk for various activation functions. We show that the softmax reaches the Bayes risk. We combine analytical arguments with numerical evidence to support the claim that this approach also captures the performance that is efficiently achievable by running one-pass SGD, which directly minimizes the population risk from random initialization. All proofs of the results in this section are deferred to Appendix A. A summary of the terminology and the notations for the risks is given in Table 1 left.
@@ 4.1 Bayes risk and optimality of softmax
We turn our attention to the best reachable theoretical risk over the class of estimators $\mathcal{F}_{\sigma}$ when $\sigma$ is the softmax function. Surprisingly, this estimator reaches the Bayes error as shown next.
@@
Note that the requirement on $g_{\nu}$ defining the distribution of $X$ holds in particular for the spiked-SLR and the max-SLR. In consequence, the softmax architectures $\mathcal{F}_{\rm softmax}$ reach the Bayes error in both settings, and furthermore the proof reveals that the minimum is reached for $k=c_{\nu}k^{*}$ and $v=v^{*}$, corresponding to recovery of the hidden directions. In statistical physics terminology, the softmax attention is said to satisfy the Nishimori condition, as detailed in the proof. This is to be contrasted with the performance of other activation functions, which is characterized next.
@@ 4.2 Expression of the population risk for arbitrary activation functions
The population risk of the attention $(k,v)\mapsto\mathcal{E}_{\sigma}(k,v)$ can be reparametrized as a function of the following 7 variables, referred to as order parameters:
@@
The first two order parameters, $m_{kk^{*}}$ and $m_{vv^{*}}$, quantify the recovery of the hidden directions, while $q_{kk}$ and $q_{vv}$ are the squared norm of the parameters $k$ and $v$. Finally, the cross-correlation terms $m_{kv^{*}}$, $m_{vk^{*}}$, and $q_{vk}$ are nuisance parameters. We let $\tilde{\mathcal{E}}_{\sigma}:\mathbb{R}^{7}\to\mathbb{R}$ be the reparametrized risk depending on the order parameters.
@@
The intuition behind this restriction is that, in the SLR, the signals $k^{*}$ and $v^{*}$ associated to the keys and the values are drawn independent ; and to obtain good performances, the keys and the values should not mix and should focus on $k^{*}$ or on $v^{*}$ separately. This simplification is supported by the following observations pertaining to the landscape of the population risk $\mathcal{E}_{\sigma}$. In practice, we do not have access to minimizers of the risk, but can instead optimize parameters $k$ and $v$ by one-pass SGD on $\mathcal{E}_{\sigma}$, which corresponds on average to running gradient descent (GD) on the reparametrized risk $\tilde{\mathcal{E}}_{\sigma}$. In the high-dimensional limit $D\to\infty$, random Gaussian initialization of the parameters lands on $\mathcal{M}$ because the cross-correlations vanish. Then, the manifold is invariant by GD, in the sense that iterates initialized on the manifold remain on it (see Appendix A.2 for details on this statement and the following). At finite $D$, random initialization lands in a neighborhood of $\mathcal{M}$. In this case, we numerically check that $\mathcal{M}$ is stable in the sense that parameters stay close throughout GD. For linear attention, we also show that the risk has only one (local) minimizer on $\mathcal{M}$, which is also a local minimizer of the risk over the whole parameter space, while for softmax attention we know that a global minimizer is on $\mathcal{M}$. Taken together, these facts support that GD on $\tilde{\mathcal{E}}_{\sigma}$ and thus one-pass SGD on $\mathcal{E}_{\sigma}$ converges to a minimizer of the risk on $\mathcal{M}$. A formal proof is delicate, as the landscape features other local minima outside of the manifold, as shown in Appendix A.2.6. This is left for future work.
With a small abuse of notation, we let $\tilde{\mathcal{E}}_{\sigma}(m_{kk^{*}},m_{vv^{*}},R_{kk},R_{vv})$ be the reparametrized risk restricted to $\mathcal{M}$, with $R_{kk}^{2}=q_{kk}-m_{kk^{*}}^{2}$ and $R_{vv}^{2}=q_{vv}-m_{vv^{*}}^{2}$ the orthogonal components. It turns out that this object is amenable to insightful theoretical analysis. We first write it explicitly.
@@
We next state that on $\mathcal{M}$ the risk of the linear attention admits a unique minimum, which together with Corollary 4.1 guarantees that GD initialized on $\mathcal{M}$ will converge to it.
Corollary 4.2.
The population risk $\mathcal{E}_{\sigma}$ of the linear attention $\sigma(\chi)_{\ell}=1+\chi_{\ell}$ admits a unique minimizer over $\mathcal{M}$. It is moreover a (local) minimizer over the whole space $(\mathbb{R})^{2}$.
@@ 4.3 Comparison between linear and softmax attentions
Letting $\mathsf{E}_{\sigma}=\min_{\mathcal{M}}\mathcal{E}_{\sigma}=\min_{\mathbb{R}^{4}}\tilde{\mathcal{E}}_{\sigma}$ be the minimum of the population risk, we now compare the minimal population risk $\mathsf{E}_{\mathrm{lin}}$ for linear attention $\sigma(\chi)_{\ell}=1+\chi_{\ell}$ to the one for softmax $\mathsf{E}_{\mathrm{softmax}}=\mathcal{E}_{\mathrm{Bayes}}$, encompassing asymptotic regimes of strong signals $\nu\to\infty$ or long sequences $L\to\infty$.
@@
$\mathsf{E}_{\mathrm{lin}}=1-\frac{L+\nu(L-1)}{L^{2}+\nu(L-1)}{\sim}\frac{L}{L-1}\frac{1}{\nu}\quad\text{while }\quad\mathsf{E}_{\mathrm{softmax}}=e^{-c_{L}\nu+o(\nu)}\quad\text{as }\nu\to+\infty$ (12)
@@
Figure 2: Minimal population risk $\mathsf{E}_{\sigma}$ over $\mathcal{F}_{\sigma}$ for different attention activations $\sigma$ (colors), compared to the Bayes risk $\mathcal{E}_{\mathrm{Bayes}}$ (5) (black), for the two tasks spiked-SLR (top) and max-SLR (bottom). Softmax is the only one achieving the Bayes risk. The markers on the lines are for readability only. Population risks are computed via numerical optimization of (9). In all cases, we found that $R_{kk}=R_{vv}=0$ was optimal, i.e. $k$ exactly aligns with $k^{*}$ and $v$ with $v^{*}$.
@@ 5 Performance of the attention at finite sample complexity
So far, our analysis has focused on the properties of the minimizer of the population risk. In practice, however, we only have access to a finite dataset $\mathcal{D}=\left\{(X_{\mu},y_{\mu})\in\mathbb{R}^{L(\mu)\times(D+1)}:\mu\in[N]\right\}$, and must therefore rely on an empirical estimator. In this section, we consider the regularized empirical risk (or training loss) defined for some regularization strengths $r_{k}\geq 0$ and $r_{v}\geq 0$ by
@@
We investigate the high-dimensional proportional limit where $N,D\to\infty$ at finite $\nicefrac{{N}}{{D}}\to\alpha=\Theta(1)$. In this part we take $\alpha$ fixed w.r.t. $N$ and $D$; letting $\alpha\to\infty$ one would recover the population risk discussed in Section 4. For simplicity, we consider the noiseless task $\Delta=0$ in the following. A summary of the terminology and the notations for the empirical risks is given in Table 1 right.
While in Section 4 the benchmark was the Bayes risk, at finite sample complexity it is natural to consider instead the best achievable test risk conditionally on a finite batch of data $\mathcal{D}$, also known as the Bayes-optimal risk (or Bayes-optimal error). In the proportional high-dimensional limit, the Bayes-optimal error for the SLR task can be computed by extending the results of Troiani et al. (2025) to random sequence lengths $L\sim P_{L}$ and arbitrary $g_{\nu}$. Although this result is of independent interest, for conciseness we refer the reader to Appendix C for the details. The outcome is that the Bayes-optimal error presents a rich phenomenology, with a hard phase where the best-known first-order method fails to achieve the information-theoretical performance, see Appendix B for a full discussion.
@@
As in the population result of Proposition 4.3, the high-dimensional analysis shows that the risk depends only on a few order parameters. To give some intuition, they can be interpreted in the following way. $m_{k}$ and $m_{v}$ are the alignments of $k$ and $v$ with $k^{*}$ and $v^{*}$, respectively. $Q_{k}$ and $Q_{v}$ are the squared norms of $k$ and $v$. $V_{k}$ and $V_{v}$ are related to the local curvature of the empirical risk $\mathcal{L}$; they encode the fluctuations due to the finite sampling ratio and when $\alpha\to\infty$ we have $V_{k},V_{v}\to 0$. Last, in the self-consistent equations, the potential $\psi_{\mathrm{out}}$ plays the role of an effective low-dimensional empirical risk. Although the resulting expressions are cumbersome, it is important to emphasize that these quantities are deterministic and independent of the diverging dimensions $N$ and $D$, in contrast to the original risk (15). Thus the result provides an implementable formula for the performance of ERMs. In practice, the minimum over the set $\mathcal{S}$ is computed by running the fixed point method from several initializations, typically random (so-called uninformed) and informed ones corresponding to (partial) alignment of $k$ and $v$ with $k^{*}$ and $v^{*}$.
@@
Figure 3: Minimal test risk of the attention (linear vs. softmax) across different tasks and signal strengths $\nu$, for $L=3$. Linear attention is shown in red and softmax in blue. Solid lines indicate $\mathsf{E}_{\sigma}(\alpha)$ at finite $\alpha$ (Result 5.1), while markers represent the test risk of an ERM obtained via a local optimization method with $\sqrt{ND}=10^{4}$. The regularizations $r_{k}$ and $r_{v}$ are tuned by grid search to minimize the test risk, as detailed in Appendix D. Dotted and dashed lines correspond to the value of $\mathsf{E}_{\sigma}$ in the infinite-$\alpha$ limit (see closed-formed formulas in Proposition 4.1 for softmax and Appendix A.2.7 for linear). The Bayes-optimal risk is shown in black (see Appendix B for a discussion on its discontinuity). Appendices D-E include more experimental details and results.
@@
We conclude by a few computational remarks. Because the empirical risk (15) is non-convex, it is not guaranteed a priori that optimization algorithms can find global minima. To assess the effect of non-convexity, we rely on numerical simulations shown in Fig. 3. More precisely, we compare the prediction of Result 5.1 (here, uninformed and informed initializations give the same result) to the outcome of running a local optimization algorithm (specifically, a quasi-Newton method) on the risk (15), starting from a random initialization, for finite but large $N$ and $D$. We first note that the agreement between both is excellent. This suggests that potential bad local minima in the optimization landscape are avoided, at least for the appropriate regularization strength. Analyzing this landscape is an interesting open question. Furthermore, the softmax attention has lower error than the linear attention across all the tested hyperparameters, which shows that the benefits of softmax extend beyond the population risk analysis of Section 4 to statistical and computational advantages on the empirical risk. We finally note that in this case the softmax is no longer Bayes-optimal but the gap to the Bayes-optimal risk closes as $\alpha$ grows, as expected from our analysis in Section 4.